A man is known to speak truth 3 out of 4 times.He throws adie and reports it is a six. Find the probability that it is actually a six.  ANS:3/8


Grade:12

1 Answers

sneha reddy
11 years ago

1. we can get a six if a man says truth =3/4*1/6

or when he lies and it is not a six=1/4*5/6

so by bayes theorem=3/4*1/6 div by 1/4*5/6+3/4*1/6=3/8